# MotionCharacter: Identity-Preserving and Motion Controllable Human Video Generation

### v2

@@ 3.3 Motion Control Enhancement
3.4 Training Paradigm
Human-Motion Dataset. To support high-quality video generation, we constructed a diverse dataset named Human-Motion, comprising 106,292 video clips from various public and private sources. This collection includes VFHQ [40] (1,843 clips), CelebV-Text [43] (52,072 clips), CelebV-HQ [46] (31,004 clips), AAHQ [21] (17,619 clips), and a private dataset, Sing Videos (3,752 clips). Each clip in the Human-Motion dataset was rigorously filtered and re-annotated to ensure high-quality identity and motion information across diverse video formats, resolutions, and styles.
To enrich the dataset with motion-related information, we used MiniGPT [45] to automatically generate two types of captions for all videos: overall descriptions and action phrases. The overall descriptions provide a general summary of the video’s content, while the action phrases offer specific annotations of facial and body movements present in the clips. These captions serve as the primary text description $\mathcal{P}$ and action phrase $\mathcal{A}$ in our framework.
Image-Video Training Strategy. To improve the model’s generalization across different visual styles, we combined image and video data in training. While realistic videos effectively capture human portraits, they struggle with stylized and artistic content, such as anime. To bridge this gap, we incorporated around 17,619 styled portrait images as static 16-frame videos by replicating each image to simulate a motionless sequence with a motion intensity of 0. This approach addresses the challenge of generalizing to stylized portraits by expanding the model’s exposure to a wider spectrum of visual characteristics, including variations in texture, color, and artistic exaggeration common in non-realistic styles. By training on both static styled images and dynamic realistic videos, the model learns to preserve identity traits across photorealistic and stylized visual representations, improving its ability to generalize across different visual styles.
@@ 3.4 Training Paradigm
$\mathcal{L}_{\text{total}}=\mathcal{L}_{R}+\beta\cdot\mathcal{L}_{id},$ (11)
where $\mathcal{L}_{R}$ guides the model to capture dynamic motion in high-activity regions, and $\mathcal{L}_{id}$ ensures identity consistency across frames. The hyperparameter $\beta$ balances the influence of identity preservation against motion fidelity.
4 Experiments
4.1 Experiment Setup
Figure 3: Qualitative Comparison. Comparison of our method with other approaches across diverse prompts and unseen reference images, encompassing various identities (male, female, celebrity, non-celebrity). Each column represents a unique identity and action phrase, with motion intensity fixed at 20 for clarity. “null” indicates a blank action phrase. Key prompt elements are highlighted in underline to emphasize specific actions or descriptors. For other methods, the action phrase and motion intensity are incorporated with the prompt to guide generation. To simplify notation, we abbreviated method names on the far left by omitting the common “FaceID” field, resulting in labels like IPA-Portrait.
@@ 4.1 Experiment Setup
Implementation Details. Our implementation is built upon a large-scale pre-trained Video Diffusion Model (VDM) [9]. All experiments were conducted using 8 NVIDIA A100 GPUs (80GB), with the training process taking approximately 24 hours. The batch size was set to 2 for each GPU. The training data consisted of diverse video clips, which were preprocessed to a resolution of $512\times 512$ pixels, with 16 frames sampled per video at a frame rate of 4 frames per second. Data augmentation included random horizontal flipping, resizing, and center cropping to maintain consistent input dimensions. Moreover, text and image dropout rates were set to 0.05, with a 50% probability of dropping the CLIP embeddings $E_{clip}$. We used the AdamW [22] optimizer with a learning rate of $1\times 10^{-5}$ and trained the model for 12,000 steps. For the validation, 16-frame video sequences were generated at a $512\times 512$ resolution, applying a guidance scale of 8.0 and 30 steps.
@@
Datasets. For training, we constructed a dataset of 106,292 video clips from various sources. We detail the composition and annotation process in Sec. 3.4. For evaluation, we used the open-source Unsplash-50 test set [8], which includes 50 portrait images sourced from the Unsplash website. For each reference image, we generated 140 prompts using GPT-4 [1], yielding a total of 7,000 prompt-image pairs for calculating evaluation metrics.
@@
Evaluation Metrics. We assess the quality and consistency of generated videos using six key metrics. The Dover Score [37] assesses overall video quality, considering technical and aesthetic factors. Motion Smoothness [17] evaluates the continuity of movement between frames, while Dynamic Degree [17] indicates the extent of motion diversity in the video. CLIP-I [12, 39] measures visual similarity to the reference using the CLIP encoder [25], and CLIP-T [12, 39] evaluates the alignment between the video content and the text description. To assess identity preservation, we calculate Face Similarity [39], which measures the resemblance between the facial features in the reference image and the generated video.
4.2 Comparison with Baselines
@@
We employ four well-known methods in ID-preserving generation task for comparison, i.e., IPA-PlusFace [42], IPA-FaceID-Portrait [42], IPA-FaceID-PlusV2 [42] and ID-Animator [11]. They all adopt AnimateDiff [9] as the base text-to-video generation model.
Qualitative Comparisons. We choose six different individuals covering celebrities and common individuals, and produce corresponding text prompts using GPT-4 [1] to present a fair comparison. As illustrated in Fig. 3, the upper videos are generated with “null” action phrase while the bottom videos are generated with “smiling”. Obviously, our proposed MotionCharacter yields superior results in identity consistency and motion control ability. The frames generated by IPA-FaceID-Portrait are less similar to the reference image, and IPA-PlusFace and IPA-FaceID-PlusV2 cannot maintain the consistency in generated video frames. Compared to ID-Animator, our method shows better capacity to align the prompt and identity information, while ID-Animator fails to present “glasses” in the generated videos.
Quantitative Comparisons. We also perform quantitative comparison with these ID-persevering methods across key metrics, including Dover Score, Motion Smoothness, Dynamic Degree, CLIP-I, and CLIP-T. As the results shown in Table 1, IPA-PlusFace performs well in Motion Smoothness but shows limited motion diversity. While IPA-FaceID-PlusV2 achieves the highest Face Similarity score, it compromises motion diversity. ID-Animator shows limitations in alignment between the video content and the text description. In contrast, our method maintains high identity consistency while producing more dynamic motion.
User Study. We recognize that the CLIP score [12, 39] may not consistently align with human perception [24, 34]. To address this, we conducted a user study to compare the quality of videos generated by our model against baselines. Participants viewed clips generated with varying action phrases and intensity levels, rating the top method on identity consistency, motion controllability, and overall video quality. We then calculated the percentage distribution across all methods. As shown in Fig. 4, our method was consistently preferred by users. This feedback aligns with our quantitative results, demonstrating its superior balance of identity preservation, motion control, and visual fidelity compared to baselines.
@@
Figure 5: Ablation study on the effects of Region-Aware Loss $\mathcal{L}_{R}$ and ID-Consistency Loss $\mathcal{L}_{id}$.
4.3 Ablation Study
Region-Aware Loss. Fig. 5 illustrates the impact of Region-Aware Loss in addressing motion blur and object distortion. In the ”Vanilla” model (top row) without Region-Aware Loss, high-motion regions like the lower corner of the glasses frame show noticeable distortion, compromising structural integrity. By adding Region-Aware Loss (middle row), the model better preserves structure in dynamic areas, resulting in clearer and more stable motion.
ID-Consistency Loss. The bottom row of Fig. 5 demonstrates the effect of adding ID-Consistency Loss on top of Region-Aware Loss. This combination significantly enhances the retention of identity-specific features, increasing facial similarity to the reference ID image. Furthermore, it better preserves facial detail characteristics (e.g., skin tone), resulting in improved overall facial integrity.
Figure 6: Ablation study on the effects of Motion Control Module.
Motion Control Module. To evaluate the effectiveness of the Motion Control Module (MCM), we conducted an ablation study, holding all training and inference parameters constant while isolating the MCM’s contribution. Specifically, we trained models both with and without the MCM, leaving all other modules unchanged. As shown in Fig. 6, the model with the MCM generates videos that follow action prompts more precisely, with enhanced clarity and sharper motion details. This improvement demonstrates that the MCM significantly boosts the model’s ability to capture nuanced action dynamics, achieving smoother transitions and more lifelike motion fidelity in the generated outputs.
5 Limitations
While our framework achieves significant performance in identity-consistent and controllable Text-to-Video (T2V) generation, it has limitations in handling highly complex or intricate motion sequences, where fine-grained motion dynamics may not be captured effectively. Additionally, the framework’s performance is inherently dependent on the capabilities of the underlying T2V base model, which can limit the quality of generated videos. As T2V base models advance, our approach is designed with potential adaptability in mind; future iterations may leverage more powerful video foundation models, such as CogVideo-X[41], to enhance generalization and video fidelity in increasingly demanding scenarios.
6 Conclusions
In this paper, we propose a framework named MotionCharacter for human video generation that emphasizes identity consistency and precise motion control. We introduce the ID-Preserving Module, which ensures stable identity representation across frames, enhancing identity fidelity in the generated video. Additionally, we present the Motion Control Module, allowing nuanced adjustments of action phrases and motion intensity for fine-grained motion dynamics. To further improve model’s performance, we leverage the region-aware loss, which reinforces fidelity in high-motion regions and achieves improved identity coherence. The Human-Motion dataset with detailed motion annotations enhances our model’s adaptability to diverse prompts. Comprehensive evaluations confirm the effectiveness of our method in generating lifelike, personalized videos that accurately capture the specified actions and motion intensities.
A Additional Dataset Analysis
@@
To build the Human-Motion dataset, a multi-step pipeline was developed to ensure the collection of high-quality video clips. The detailed process is illustrated in Fig. I.
A.1 Video Sources
@@
Our data sources comprise video clips from diverse origins, including VFHQ [40], CelebV-Text [43], CelebV-HQ [46], AAHQ [21], and a private dataset, Sing Videos. Each clip was carefully filtered and re-annotated to ensure high-quality identity and motion information across various formats, resolutions, and styles.
@@ A.2 Filtering Process
  • Visual Quality Check: We used CLIP Image Quality Assessment [31] (CLIP-IQA) to evaluate visual quality by sampling one frame per clip, discarding videos with frames of low quality.
@@
  • Text Overlay Detection: EasyOCR [18] was used to detect excessive subtitles or text overlays, filtering out obstructed frames.
@@
A.3 Captioning
To enrich motion-related data, we utilized MiniGPT [45] to automatically generate two types of captions for each video:
@@ A.3 Captioning
A.4 Optical Flow Estimation
Optical flow estimation for video is performed using the RAFT [30] model on consecutive frames to compute motion information. The RAFT model calculates the optical flow field, representing pixel displacements between frames. In this study, we use the extracted optical flow to obtain motion information for each video segment, enabling accurate motion modeling and control. These optical flows are used to compute motion intensity during the training phase and are also utilized in optimizing the loss function.
A.5 Motion Intensity Resampling
@@
A.6 The Human-Motion Dataset
The Human-Motion dataset consists of 106,292 video clips sourced from various datasets, including VFHQ [40] (1,843 clips), CelebV-Text [43] (52,072 clips), CelebV-HQ [46] (31,004 clips), AAHQ [21] (17,619 clips), and a private dataset, Human Videos (3,752 clips). Each clip was rigorously filtered and re-annotated to ensure high-quality identity and motion information across diverse formats, resolutions, and styles.

### v3

@@ Introduction
Training Paradigm
Our training paradigm is tailored for learning fine-grained, disentangled control. It comprises three integral parts: (1) a dataset with dual-track annotations for motion type and intensity, (2) a mixed image-video strategy for robust generalization and zero-motion calibration, and (3) a loss function that synergistically optimizes for both identity stability and motion fidelity.
Human-Motion Dataset. While foundational, current human video datasets like CelebV-HQ (zhu2022celebv) and CelebV-Text (yu2023celebv) lack the fine-grained motion annotations and quality controls needed for training controllable models. Their coarse emotional categories and data artifacts limit the learning of nuanced dynamics. To enable disentangled motion generation, we introduce Human-Motion, comprising 106,292 video clips from various public and private sources. This collection includes VFHQ (xie2022vfhq) (1,843 clips), CelebV-Text (yu2023celebv) (52,072 clips), CelebV-HQ (zhu2022celebv) (31,004 clips), AAHQ (liu2021blendgan) (17,619 clips), and a private dataset (3,752 clips). Each clip in the Human-Motion dataset was rigorously filtered and re-annotated to ensure high-quality identity and motion information across diverse video formats, resolutions, and styles.
Human-Motion is distinguished by its dual-track annotation schema, which provides parallel labels for both motion semantics and intensity. Specifically, to enrich the dataset with motion-related information, we used LMM (zhu2023minigpt) to automatically generate two types of captions for all videos: overall descriptions and action phrases. The overall descriptions provide a general summary of the video’s content, while the action phrases offer specific annotations of facial and body movements present in the clips. These captions serve as the primary text description $\mathcal{P}$ and action phrase $\mathcal{A}$ in our framework.
Image-Video Training Strategy. To improve the model’s generalization across different visual styles, we combined image and video data in training. While realistic videos effectively capture human portraits, they struggle with stylized and artistic content, such as anime. To bridge this gap, we incorporated around 17,619 styled portrait images as static 16-frame videos by replicating each image to simulate a motionless sequence with a motion intensity of 0. It provides a crucial “zero-intensity calibration” for our motion control module. Training on these static sequences enables the model to learn a robust “motionless” state representation, which anchors the zero-point of the continuous motion intensity spectrum. This calibration contributes to generating smooth, predictable transitions from complete stillness to dynamic action. Moreover, this approach addresses the challenge of generalizing to stylized portraits by expanding the model’s exposure to a wider spectrum of visual characteristics, including variations in texture, color, and artistic exaggeration common in non-realistic styles. By training on both static styled images and dynamic realistic videos, the model learns to preserve identity traits across photorealistic and stylized visual representations, improving its ability to generalize across different visual styles.
@@
$\mathcal{L}_{\text{total}}=\mathcal{L}_{R}+\lambda_{id}\cdot\mathcal{L}_{id},$ (11)
where $\mathcal{L}_{R}$ and $\mathcal{L}_{id}$ synergistically guide the optimization. $\mathcal{L}_{R}$ ensures motion fidelity by focusing on dynamic regions, while $\mathcal{L}_{id}$ preserves identity consistency across the sequence. The hyperparameter $\lambda_{id}$ mediates the trade-off between these two critical objectives.
Experiments
Experiment Setup
@@
Implementation Details. Our implementation is built upon a large-scale pre-trained Video Diffusion Model (VDM) (guo2023animatediff). All experiments were conducted using 8 NVIDIA A100 GPUs (80GB), with the training process taking approximately 24 hours. The batch size was set to 2 for each GPU. The training data consisted of diverse high-quality video clips, which were preprocessed to a resolution of $512\times 512$ pixels, with 16 frames sampled per video at a frame rate of 4 frames per second. Data augmentation included random horizontal flipping, resizing, and center cropping to maintain consistent input dimensions. Moreover, text and image dropout rates were set to 0.05, with a 50% probability of dropping the CLIP embeddings $E_{clip}$. We used the AdamW (loshchilov2017decoupled) optimizer with a learning rate of $1\times 10^{-5}$ and trained the model for 12,000 total steps. For the validation, 16-frame video sequences were generated at a $512\times 512$ resolution, applying a standard guidance scale of 8.0 and 30 steps.
@@
Datasets. For training, we constructed a dataset of 106,292 video clips from various sources. We detail the composition and annotation process in the Training Paradigm subsection. For evaluation, we benchmark our method on the Unsplash-50 test set (gal2024lcm), a standard benchmark for ID fidelity containing 50 diverse portraits. To rigorously test performance, we paired each portrait with 140 distinct prompts generated via GPT-4 (achiam2023gpt), yielding a challenging evaluation suite of 7,000 pairs.
@@
Evaluation Metrics. We assess the quality and consistency of generated videos using six key metrics. The Dover Score (wu2023dover) assesses overall video quality, considering technical and aesthetic factors. Motion Smoothness (huang2023vbench) evaluates the continuity of movement between frames, while Dynamic Degree (huang2023vbench) indicates the extent of motion diversity in the video. CLIP-I (hessel2021clipscore; xiao2024fastcomposer) measures visual similarity to the reference using the CLIP encoder (radford2021learning), and CLIP-T (hessel2021clipscore; xiao2024fastcomposer) evaluates the alignment between the video content and the text description. To assess identity preservation, we calculate Face Similarity (xiao2024fastcomposer), which measures the resemblance between the facial features in the reference image and the generated video.
Comparison with Baselines
@@
We employ four well-known methods in ID-preserving generation task for comparison, i.e., IPA-PlusFace (ye2023ip), IPA-FaceID-Portrait (ye2023ip), IPA-FaceID-PlusV2 (ye2023ip) and ID-Animator (he2024id). They all adopt AnimateDiff (guo2023animatediff) as the base Text-to-Video generation model.
Figure 3: Qualitative Comparison. Comparison of our method with other approaches across diverse prompts and unseen reference images, encompassing various identities (male, female, celebrity, non-celebrity). Each column represents a unique identity and action phrase, with motion intensity fixed at 20 for clarity. “null” indicates a blank action phrase. Key prompt elements are highlighted in underline to emphasize specific actions or descriptors. For other methods, the action phrase and motion intensity are incorporated with the prompt to guide generation.
Qualitative Comparisons. We conduct qualitative comparisons using six diverse identities and prompts generated by GPT-4 (achiam2023gpt). As shown in Fig. 3, we test a static scenario (null action phrase) to assess baseline ID fidelity and a dynamic scenario (“smiling”) to evaluate motion control. The visual results underscore the superiority of MotionCharacter. Methods like IPA-FaceID-Portrait exhibit noticeable identity degradation, while others like IPA-PlusFace and IPA-FaceID-PlusV2 suffer from temporal flickering during motion. This indicates their ID representations are not robust enough to withstand dynamic deformations. Furthermore, even the highly competitive ID-Animator fails to bind attributes correctly, omitting “glasses” from the generated character. This suggests a critical entanglement between its motion and appearance generation modules. In contrast, our method, powered by a dedicated ID-Preserving module and a disentangled motion architecture, successfully maintains identity and all specified attributes while executing the action.
[…]
  IPA-PlusFace    0.797 0.985 0.325 0.587 0.218 0.480
IPA-FaceID-Portrait    0.849 0.984 0.191 0.545 0.223 0.531
IPA-FaceID-PlusV2    0.813 0.987 0.085 0.575 0.217 0.617
ID-Animator    0.857 0.979 0.433 0.607 0.204 0.546
Ours  0.869 0.998 0.449 0.633 0.227 0.609
Table 1: Quantitative comparison with state-of-the-art methods. Higher values ($\uparrow$) indicate better performance. Bold and underlined numbers denote the best and second-best results, respectively. All methods use an empty action phrase with motion intensity set to 20 for fair comparison.
Quantitative Comparisons. The quantitative results in Table 1 further substantiate our findings and reveal a critical challenge in existing methods: the inherent trade-off between identity preservation and motion dynamics. Specifically, IPA-FaceID-PlusV2 achieves the highest Face Similarity (0.617) but at the cost of producing nearly static videos, evidenced by its extremely low Dynamic Degree (0.085). Conversely, ID-Animator attains a high Dynamic Degree (0.433) but with significantly compromised identity fidelity and weaker text alignment (CLIP-T of 0.204), which confirms our qualitative observations. This demonstrates that current methods are forced to sacrifice either motion or identity. In contrast, MotionCharacter breaks this trade-off. Our method achieves state-of-the-art scores in five out of six metrics, including overall quality (Dover), motion attributes, and content consistency (CLIP-I, CLIP-T). Crucially, we attain a Face Similarity score (0.609) that is highly competitive with the top score, while simultaneously achieving the highest Dynamic Degree (0.449). This result is a powerful validation of our core hypothesis: by explicitly disentangling motion control from identity representation, our framework can generate highly dynamic and expressive videos without compromising identity fidelity. The negligible 1.3% difference in Face Similarity is a small price for a staggering 428% improvement in Dynamic Degree compared to the leading ID-preserving method.
@@
User Study. Recognizing that CLIP scores (hessel2021clipscore; xiao2024fastcomposer) may not fully align with human perception (molad2023dreamix; wang2023zero), we conducted a user study to validate our quantitative findings. The study involved 10 expert raters from enterprises and academic labs, each with expertise in generative models, evaluating 100 videos across three key dimensions: Identity Consistency, Motion Controllability, and Overall Video Quality. For each video, participants selected the best performing method among all approaches, with videos presented in randomized order to prevent bias. This resulted in 3,000 total ratings (10 raters $\times$ 100 videos $\times$ 3 criteria). As shown in Fig. 4, our method consistently received the highest preference across all evaluation criteria. Notably, the slight discrepancy between Face Similarity metrics and human perception of identity preservation can be attributed to varying levels of motion richness. While methods like ID-Animator and IPA achieve high similarity scores through strict identity constraints that limit facial dynamics, our approach generates more varied and expressive motions while maintaining perceptually consistent identity. This human evaluation validates our quantitative results, confirming that our decoupled control framework achieves a superior balance between identity preservation and motion expressiveness.
Ablation Study
Region-Aware Loss. As shown in Table 2, integrating only the Region-Aware Loss ($\mathcal{L}_{R}$) yields substantial improvements in motion-related metrics compared to the vanilla baseline. Specifically, it boosts the Dover Score by a significant +0.059 and increases the Dynamic Degree by +0.064, while also enhancing Motion Smoothness. This quantitatively confirms that by forcing the model to focus its objective on high-motion areas, $\mathcal{L}_{R}$ is highly effective at improving the clarity and quality of dynamic movements. Notably, its impact on Face Similarity is minimal, underscoring its specific role in refining motion rather than identity.
     Loss  Metrics
[…]
  ✗    0.801 0.978 0.355 0.599 0.219 0.484
✓    0.810 0.983 0.359 0.627 0.220 0.588
✗    0.860 0.995 0.419 0.611 0.224 0.500
✓    0.869 0.998 0.449 0.633 0.227 0.609
@@
Table 2: Ablation study of the Region-Aware Loss $\mathcal{L}_{R}$ and the ID-Consistency Loss $\mathcal{L}_{id}$.
  Module      Metrics
[…]
  ✗    0.805 0.978 0.245 0.563 0.204 0.601
✓    0.869 0.998 0.449 0.633 0.227 0.609
Table 3: Ablation study of Motion Control Module (MCM).
ID-Consistency Loss. The ID-Consistency Loss ($\mathcal{L}_{id}$) is designed to anchor the subject’s identity. When added alone, it produces a massive gain of +0.104 in Face Similarity, proving its efficacy in robust identity preservation. However, the most compelling finding is the synergistic effect when both losses are combined. The full model not only achieves the best scores across all metrics but also surpasses the performance of individual components in their respective domains. For instance, the final Face Similarity (0.609) is higher than with $\mathcal{L}_{id}$ alone (0.588), and the Dynamic Degree (0.449) is greater than with $\mathcal{L}_{R}$ alone (0.419). This powerful synergy validates our core hypothesis: a stable identity foundation provided by $\mathcal{L}_{id}$ enables $\mathcal{L}_{R}$ to sculpt more expressive and high-fidelity motion, demonstrating that both losses are critical and complementary for achieving state-of-the-art results.
Motion Control Module. Table 3 demonstrates the substantial impact of our Motion Control Module (MCM). Its introduction leads to consistent improvements across all metrics, with particularly notable gains in Dynamic Degree (83.3% increase from 0.245 to 0.449). This dramatic improvement in motion expressiveness, achieved while maintaining face similarity (0.609), validates our module’s ability to enhance motion control without compromising identity preservation.
Conclusions
In this paper, we presented MotionCharacter, a high-fidelity human video generation framework that successfully addresses the fundamental challenge of achieving fine-grained motion control while ensuring strict identity preservation. The key to our approach is a disentangled architecture where two parallel mechanisms operate. The first mechanism anchors the subject’s appearance using a dedicated ID Content Insertion Module supervised by an ID-Consistency Loss. The second enables precise control over action and intensity via the Motion Control Module, with a Region-Aware Loss ensuring clarity in dynamic areas. This entire framework is empowered by our Human-Motion dataset, whose fine-grained annotations are crucial for learning such disentangled representations. Exhaustive experiments validate that MotionCharacter not only outperforms existing methods in overall quality but also effectively breaks the critical trade-off between motion dynamics and identity fidelity.
Acknowledgments
This work was supported by the National Natural Science Foundation of China (Grant U25A20403), the Natural Science Foundation of Hubei Province of China (No. 2024AFB545). We also thank Meituan for their support.
Additional Dataset Analysis
@@ Acknowledgments
To build the Human-Motion dataset, a multi-step pipeline was developed to ensure the collection of high-quality video clips. The detailed process of building the dataset is illustrated in Fig. I.
Video Sources
@@
Our data sources comprise video clips from diverse origins, including VFHQ (xie2022vfhq), CelebV-Text (yu2023celebv), CelebV-HQ (zhu2022celebv), AAHQ (liu2021blendgan), and a private dataset, Sing Videos. Each clip was carefully filtered and re-annotated to ensure high-quality identity and motion information across various formats, resolutions, and styles.
@@
  • Visual Quality Check: We used CLIP Image Quality Assessment (wang2023exploring) (CLIP-IQA) to evaluate visual quality by sampling one frame per clip, discarding videos with frames of low quality.
@@
  • Text Overlay Detection: EasyOCR (JaidedAI_EasyOCR) was used to detect excessive subtitles or text overlays, filtering out obstructed frames.
@@
Captioning
To enrich motion-related data, we utilized MiniGPT (zhu2023minigpt) to automatically generate two types of captions for each video:
@@
Optical Flow Estimation
Optical flow estimation for video is performed using the RAFT (teed2020raft) model on consecutive frames to compute motion information. The RAFT model calculates the optical flow field, representing pixel displacements between frames. In this study, we use the extracted optical flow to obtain motion information for each video segment, enabling accurate motion modeling and control. These optical flows are used to compute motion intensity during the training phase and are also utilized in optimizing the loss function.
Motion Intensity Resampling
@@
The Human-Motion Dataset
The Human-Motion dataset consists of 106,292 video clips sourced from various datasets, including VFHQ (xie2022vfhq) (1,843 clips), CelebV-Text (yu2023celebv) (52,072 clips), CelebV-HQ (zhu2022celebv) (31,004 clips), AAHQ (liu2021blendgan) (17,619 clips), and a private dataset, Sing Videos (3,752 clips). Each clip was rigorously filtered and re-annotated to ensure high-quality identity and motion information across diverse formats, resolutions, and styles.
Additional Visual Results
Action Phrase Control
To evaluate our model’s responsiveness to different action instructions, we conduct controlled experiments by varying only the action phrases while keeping the reference image, text prompt, and motion intensity constant. As shown in Fig. II, our model accurately interprets and executes diverse action phrases such as “turn head”, “smile”, and “talk, hold a microphone”.
Figure II: Action phrase control demonstration. Our model generates corresponding motions for different action phrases (e.g., “turn head”, “smile”, “null”, “talk, hold a microphone”) while maintaining consistent identity and context. The results showcase MotionCharacter’s ability to handle both simple and compound action instructions.
Figure III: Motion intensity control demonstration. With fixed reference image and action phrase (“open mouth”), we show generated results at different intensity levels. Lower values (5) produce subtle movements, while higher values (20) generate more pronounced actions.
The results demonstrate two key capabilities: (1) precise action execution - each phrase generates distinct and appropriate movements, from subtle facial expressions to more complex combined actions; and (2) identity preservation - the subject’s core features remain stable across different motions, validating our framework’s ability to disentangle identity and motion attributes.
Motion Intensity Control
We analyze our model’s ability to modulate motion intensity by fixing all parameters except the intensity value. Fig. III shows the results of varying intensity levels (5, 10, 20) for the same action phrase “open mouth”.
The visualizations reveal smooth gradation of motion magnitude: lower intensity values (5) produce subtle movements while maintaining high temporal coherence, whereas higher values (20) generate more pronounced actions without compromising motion quality or identity consistency.
Additional Implementation Details
Our implementation is built upon a large-scale pre-trained Video Diffusion Model (VDM) (guo2023animatediff). All experiments were conducted using 8 NVIDIA A100 GPUs (80GB), with the training process taking approximately 24 hours. The re-annotation of the video and image datasets used in our experiments was performed in parallel on 32 NVIDIA V100 GPUs (32GB). The batch size was set to 2, with gradient accumulation over 20 steps to effectively simulate a larger batch size, allowing for efficient GPU memory usage and stable training. The training data consisted of diverse video clips, which were preprocessed to a resolution of $512\times 512$ pixels, with 16 frames sampled per video at a frame rate of 4 frames per second. Face detection and alignment were performed using the InsightFace library (deng2019arcface) to ensure well-aligned facial regions for training. Data augmentation included random horizontal flipping, resizing, and center cropping to maintain consistent input dimensions. Furthermore, text and image dropout rates were set to 0.05, with a 50% probability of dropping the CLIP embeddings $E_{clip}$. These strategies enhanced the model’s robustness by simulating various levels of input occlusion and missing information. To optimize training efficiency, we applied mixed precision using fp16 to reduce memory consumption, along with gradient checkpointing. Additionally, xFormers memory-efficient attention (xFormers2022) was activated to further optimize memory usage. We used the AdamW (loshchilov2017decoupled) optimizer with a learning rate of $1\times 10^{-5}$. Our model was trained for a total of 12,000 steps. During training, only the ID-Preserving Adapter and Motion Control Module are optimized, with all other model components kept frozen.
Limitations
While our framework achieves significant performance in identity-consistent and controllable Text-to-Video (T2V) generation, it has limitations in handling highly complex or intricate motion sequences, where fine-grained motion dynamics may not be captured effectively. Additionally, the framework’s performance is inherently dependent on the capabilities of the underlying T2V base model, which can limit the quality of generated videos. As T2V base models advance, our approach is designed with potential adaptability in mind; future iterations may leverage more powerful video foundation models, such as Wan series (wan2025wan), to enhance generalization and video fidelity in increasingly demanding scenarios.